### v1

@@ 1 Introduction
Cyber-Physical Systems (CPS) continuously generate large-scale and complex time series data across industrial domains such as manufacturing, process engineering, and energy distribution, which enable vital tasks including Prognostics and Health Management (PHM) [1]. Deep Learning (DL) has proven effective in modeling complex data, yet despite its predictive capabilities the adoption of DL in industrial CPS remains limited due to concerns over robustness [2]. Robustness, in this context, refers to the model’s ability to maintain reliable performance when encountering erroneous or unforeseen inputs [3]. Regulatory developments, such as the European Union’s AI Act, have further highlighted the importance of robust DL systems, urging systematic methodologies for robustness assessment [4]. However, consensus on practical robustness testing methodologies remains elusive [5]. While formal verification methods and adversarial robustness studies have been prominent in the literature, their practical relevance to industrial scenarios is limited due to unrealistic assumptions [6]. Recent research emphasizes evaluating model robustness against realistic data quality issues and distributional shifts commonly encountered in real-world settings [7, 8, 9, 10]. However, most research on robustness for CPS either defines robustness theoretically but lacks an extensive evaluation on real CPS data or relies on evaluations on synthetic datasets while missing a comprehensive definition of robustness.
@@ 2 Related Work
Robustness in DL: Robustness of a DL model is generally described as the ability "to maintain its level of performance under any circumstances" [3], or its capability "to cope with erroneous, noisy, unknown, and adversarial input data" [11]. It is closely related to other concepts such as reliability [12], resilience [13], safety [6] and stability [14]. DL models are known to be highly sensitive to small input perturbations, as first demonstrated by Szegedy et al. [15]. Consequently, much research has focused on adversarial robustness, evaluating model resilience to these small, intentionally crafted perturbations [16, 17], leading to formal guarantees and metrics to measure this type of robustness [18, 19]. However, adversarial perturbations rarely occur naturally, thus recent research emphasizes measuring robustness to broader, more realistic distributional shifts [20, 21]. Traditional adversarial robustness metrics often fail to sufficiently capture these distributional shifts [22, 23]. To address this limitation, recent studies propose evaluating robustness either by assessing performance across multiple distinct environments [24, 25], by applying broad yet realistic natural distribution shifts (e.g., image blur or style changes) [26], or by specifically investigating the impact of data-quality issues [27, 28].
@@
Robustness for Time Series and CPS: Robustness in time series forecasting is generally assessed by evaluating how model performance deteriorates in response to anomalies, commonly measured by increased forecasting errors [29], or by measuring the effect on the output distribution, e.g. by measuring the Wasserstein distance [30]. In CPS, research of robustness has traditionally focused on formal verification [31] or stability in control engineering [14]. Recent studies have specifically investigated the effects of data quality issues such as sensor drift, outliers and missing values on CPS data [8, 9]. Other CPS-focused robustness research has explored adversarial attacks [7], false-positive predictions under varying grades of handicap [32], and model robustness to added random noise [10]. However, most research on robustness for CPS either defines robustness theoretically but lacks an extensive evaluation on real CPS data or relies on evaluations on synthetic datasets while missing a comprehensive definition of robustness.
@@
Prognostics in CPS: In CPS contexts, PHM assesses system health and predicts future system behavior [1, 33]. Many PHM applications rely heavily on DL forecasting models [34, 35]. Commonly used architectures include recurrent neural networks (e.g., LSTM [36], GRU [37]), modular approaches like Recurrent Independent Mechanisms (RIMs) [38], Temporal Convolutional Networks (TCN) [39], attention-based models such as Transformers [40] and Informers [41], and structured state-space models like Mamba [42, 43]. Recent findings indicate that even simple feed-forward neural networks can be competitive [44]. These models represent diverse architectural approaches commonly adopted in CPS forecasting tasks. Despite their widespread application, the robustness of these architectures has primarily been evaluated using synthetic or idealized datasets, potentially masking their real-world vulnerabilities.
In summary, while extensive research has explored general and adversarial robustness in DL, and recent work has begun addressing CPS-specific robustness issues, there remains a notable gap: robustness in CPS has not been explicitly defined or comprehensively evaluated using real-world CPS datasets, which present inherent complexities. This paper addresses this gap by explicitly defining robustness tailored for CPS scenarios and evaluating it thoroughly on real-life CPS datasets, thus providing more realistic and applicable robustness insights.
@@ 3 Defining Robustness for CPS
For example, a common choice for $\mu$ is the Mean Squared Error (MSE), defined by $\text{MSE}\,(\hat{Y},Y)=\|\hat{Y}-Y\|_{2}^{2}.$ We assume that the performance measure is non-negative, with a value of 0 indicating optimal performance. Since the models are evaluated on a finite dataset, the loss $\mu$ is finite and thus integrable. If a performance measure does not naturally meet these criteria, it can be transformed via monotonic transformations.
@@ 4 Proposed Framework
To evaluate the proposed definition of robustness, we implemented a framework that simulates disturbances in CPS datasets to quantify the robustness of popular DL model architectures. We focus on a forecasting task as a representative objective, since it requires a deep understanding of the underlying system and is closely related to common industrial applications such as PHM. The framework is openly accessible, supports custom time series data in CSV or Parquet format and can easily integrate models implemented in Python. An overview of the proposed framework is shown in Figure 1.
@@
The disturbances (see Eq. (4)) simulate behavior commonly encountered in real-world CPS environments, like sensor drift, noise and irregular sampling, in gradually increasing severity. Thus, the robustness score (see Eq. (7)) can effectively measure the effect of such critical scenarios, providing a relevant quantification of model robustness. Figure 2 provides an overview of these test scenarios. For each dataset, we affect a subset of 10% of the sensors. Since the dataset is standardized, the effect of each disturbance is comparable across different datasets. Some disturbances apply solely to continuous signals (e.g., Noise), while others are specific to discrete signals or actuators (WrongDiscreteValue, OscillatingSensor). Most disturbances are not applied to the portion of the time series to be forecasted, as it would be infeasible or undesirable to predict faulty sensor values. However, MissingData requires the model to ignore the missing values and forecast using the most recent available time steps, so the section of the time series to predict changes.
@@ 7 Discussion
The disturbance robustness scores displays a large variability across some datasets, which shows the complexity in designing realistic yet challenging disturbance scenarios. This emphasizes the importance of multi-dataset, multi-scenario evaluations to reliably assess and compare model robustness.
Future work should systematically evaluate additional relevant model classes, including interpretable statistical models, hybrid statistical-DL approaches, physics-informed models, graph-based architectures, large-scale pretrained foundation models and other state-of-the-art methods. Furthermore, the disturbances could be used during training to improve model robustness.

### v2

@@ Abstract
© 2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works. DOI: XX.XXXX/XX
@@ 1 Introduction
Cyber-Physical Systems (CPS) continuously generate large-scale and complex time series data across industrial domains such as manufacturing, process engineering, and energy distribution, which enable vital tasks including Prognostics and Health Management (PHM) [1]. Deep Learning (DL) has proven effective in modeling complex data, yet despite its predictive capabilities the adoption of DL in industrial CPS remains limited due to concerns over robustness [2]. Robustness, in this context, refers to the model’s ability to maintain reliable performance when encountering erroneous or unforeseen inputs [3]. Regulatory developments, such as the European Union’s AI Act, have further highlighted the importance of robust DL systems, urging systematic methodologies for robustness assessment [4]. However, consensus on practical robustness testing methodologies remains elusive [5]. While formal verification methods and adversarial robustness studies have been prominent in the literature, their practical relevance to industrial scenarios is limited due to unrealistic assumptions [6]. Recent research emphasizes evaluating model robustness against realistic data quality issues and distributional shifts commonly encountered in real-world settings [7, 8, 9, 10]. However, existing studies stop short of combining an operational robustness metric and a large-scale evaluation on real CPS data, leaving practitioners without a practical benchmark.
@@ 2 Related Work
Robustness in DL: Robustness of a DL model is generally described as the ability "to maintain its level of performance under any circumstances" [3], or its capability "to cope with erroneous, noisy, unknown, and adversarial input data" [11]. It is closely related to other concepts such as reliability [12], resilience [13], safety [6] and stability [14]. DL models are known to be highly sensitive to small input perturbations[15]. Consequently, much research has focused on adversarial robustness, evaluating model resilience to these small, intentionally crafted perturbations [16], leading to formal guarantees and metrics to measure this type of robustness [17]. However, adversarial perturbations rarely occur naturally, thus recent research emphasizes measuring robustness to broader, more realistic distributional shifts [18, 19]. Traditional adversarial robustness metrics often fail to sufficiently capture these distributional shifts [20, 21]. To address this limitation, recent studies propose evaluating robustness either by assessing performance across multiple distinct environments [22], by applying broad yet realistic natural distribution shifts (e.g., image blur or style changes) [23], or by specifically investigating the impact of data-quality issues [24, 25].
@@
Robustness for Time Series and CPS: Robustness in time series forecasting is generally assessed by evaluating how model performance deteriorates in response to anomalies, commonly measured by increased forecasting errors [26], or by measuring the effect on the output distribution, e.g. by measuring the Wasserstein distance [27]. In CPS, research of robustness has traditionally focused on formal verification [28] or stability in control engineering [14]. Other CPS-focused robustness research has explored adversarial attacks [7], false-positive predictions under varying grades of handicap [29], and model robustness to added random noise [10]. Recent studies have specifically investigated the effects of data quality issues such as sensor drift, outliers and missing values on simulated CPS data [8, 9].
@@
Prognostics in CPS: In CPS contexts, PHM assesses system health and predicts future system behavior [1]. Many PHM applications rely on DL forecasting models, either by using them to directly predict the remaining useful lifetime (RUL) or by first forecasting health indicators of the system. In general, more accurate forecasts of health indicators are strongly associated with improved RUL prediction accuracy [30, 31]. Commonly used architectures include recurrent neural networks (e.g., LSTM [32], GRU [33]), modular approaches like Recurrent Independent Mechanisms (RIMs) [34], Temporal Convolutional Networks (TCN) [35], attention-based models such as Transformers [36] and Informers [37], and structured state-space models like Mamba [38]. Recent findings indicate that even simple feed-forward neural networks can be competitive [39]. These models represent diverse architectural approaches commonly adopted in CPS forecasting tasks. Despite their widespread application, the robustness of these architectures has primarily been evaluated using synthetic or idealized datasets, potentially masking their real-world vulnerabilities.
In summary, no prior paper combines an operational robustness score and a large-scale evaluation on real CPS data under realistic disturbances. Papers that introduce formal definitions validate them only on synthetic case studies [10, 29] or focus solely on adversarial noise rather than real-world scenarios [7]. Studies that do examine realistic data-quality issues limit their evaluations to simulations and report robustness as simple performance deltas [8, 9]. This paper closes these gaps by proposing a distribution-grounded robustness score and benchmarking nine forecasting models under representative sensor-fault scenarios across six industrial CPS data sets.
@@ 3 Defining Robustness for CPS
For example, a common choice for $\mu$ is the Mean Squared Error (MSE), defined by
$\text{MSE}\,(Y,\hat{Y})=\frac{1}{t^{\prime}n}\sum_{i=1}^{t^{\prime}}\sum_{j=1}^{n}\bigl{(}Y_{ij}-\hat{Y}_{ij}\bigr{)}^{2}.$
We assume that the performance measure is non-negative, with a value of 0 indicating optimal performance. Since the models are evaluated on a finite dataset, the loss $\mu$ is finite and thus integrable. If a performance measure does not naturally meet these criteria, it can be transformed via monotonic transformations.
@@ 4 Proposed Framework
To evaluate the proposed definition of robustness, we implemented a framework that simulates disturbances in CPS datasets to quantify the robustness of popular DL model architectures. We focus on a forecasting task as a representative objective, since it requires a deep understanding of the underlying system, which can translate into more precise RUL estimates [30, 31]. The framework is openly accessible, supports custom time series data in CSV or Parquet format and can easily integrate models implemented in Python. An overview of the proposed framework is shown in Figure 1.
@@
The disturbances (see Eq. (4)) simulate behavior commonly encountered in real-world CPS environments, like sensor drift, noise and irregular sampling, in gradually increasing severity. Thus, the robustness score (see Eq. (7)) can effectively measure the effect of such critical scenarios, providing a relevant quantification of model robustness. Ideally, the severity increment should be infinitesimally small so that the discrete summation closely approximates the integral in Eq. (6). In practice, we approximate the integral with uniform 1% severity steps, which we found to balance numerical accuracy and computational cost. Figure 2 provides an overview of these test scenarios. For each dataset, we affect a subset of 10% of the sensors. Since the dataset is standardized, the effect of each disturbance is comparable across different datasets. Some disturbances apply solely to continuous signals (e.g., Noise), while others are specific to discrete signals or actuators (WrongDiscreteValue, OscillatingSensor). Most disturbances are not applied to the portion of the time series to be forecasted, as it would be infeasible or undesirable to predict faulty sensor values. However, MissingData requires the model to ignore the missing values and forecast using the most recent available time steps, so the section of the time series to predict changes.
@@ 7 Discussion
Robustness scores varied markedly across datasets, highlighting how difficult it is to craft disturbance scenarios that are simultaneously realistic and challenging. Our present benchmark inevitably leaves several aspects open: it targets only forecasting (not anomaly detection or control) and injects sensor-level faults individually. Hence, correlated system-wide failures, actuator-side or adversarial attacks remain outside our scope, as do CPS domains such as robotics.
Future work should therefore broaden both model and scenario coverage: adding interpretable statistical or physics-informed baselines, graph-based architectures, large pretrained foundation models and incorporating disturbance-aware training or adversarial augmentation.